please help me solve this: log(x+10) - logx = 1
Answer by nerdybill(7384)

log(x+10) - logx = 1
log(x+10)/x = 1
(x+10)/x = 10^1
(x+10)/x = 10
x+10 = 10x
10 = 9x
10/9 = x